OpenStudy (anonymous):

what is 4000 divided by 80

OpenStudy (saifoo.khan):

Fifty.

OpenStudy (turingtest):

calculator saifoo...

OpenStudy (anonymous):

thx for the answer

OpenStudy (anonymous):

Well 80 divded by 400 is 5 so then add another 0

OpenStudy (saifoo.khan):

lol @turing

OpenStudy (anonymous):

You can google to find this answer. Type in "4000/80" and google will do the math for you. You've already been given the answer.

OpenStudy (anonymous):

@ saifoo.

OpenStudy (saifoo.khan):

o_O

OpenStudy (anonymous):

Grassman.. the answer is 50

OpenStudy (anonymous):

4000/80 =400/8 =50*8/8 =50

OpenStudy (saifoo.khan):

Grassman, get some sleep man!

OpenStudy (anonymous):

AHHHH, I hate you all! I'm just going to stop answering questions for a while!

OpenStudy (anonymous):

Lol, I've gotten like 4 wrong in the past 10 minutes, i think that's a new record.

OpenStudy (anonymous):

It's annoying because they've all been easy!

OpenStudy (saifoo.khan):

LOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOL

OpenStudy (turingtest):

sleep...

OpenStudy (saifoo.khan):

GN.

OpenStudy (anonymous):

It's annoying because they've all been easy!

OpenStudy (anonymous):

I never sleep! I'm a vampire, and not one of those gay vampires that live off of animal blood and go camping, I'm the real deal!

OpenStudy (anonymous):

Wait, what I didn't write aloft.........?